Please show work:

A fuel consisting of 20 mol % methane (CH4) and 80 mol % ethanol (C2H6O) is flowing at a steady rate of 12.5 mol/s through a reactor. This fuel is burned with the stoichiometric amount of dry air. What is the flow rate of the carbon dioxide (CO2) in the product stream?

Answer: [22.5 mol/s]
Answer #1

Basis 12.5 mol of fuel

0.2*12.5 = 2.5 mol of CH4

0.8*12.5 = 10 mol of C2H6O

for 1 mol of CH4 = 1 mol of CO2 = 2.5 mol of CO2

for 1 mol of C2H6O = 2 mol of CO2 = 10*2 = 20 mol of CO2

total mol = 22.5 mol of CO2

this is per unit scond so

mol of CO2 = 22.5 mol/s